OpenStudy (anonymous):

Solve 9x2 – 45x – 81 = 0.

OpenStudy (anonymous):

Allybally, you need to learn to use the quadratic formula: \[a x^2 +bx +c=0\] has solutions \[ x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

first divide thru by 9 x^2 - 5x - 9 = 0 now use quadrtic formula to find the roots

OpenStudy (anonymous):

Now just plug and chug

OpenStudy (anonymous):

Of course, you should always try to factor it first, but it seems like the questions you are getting will not factor well.

OpenStudy (anonymous):

right - it wont factor - but first dividing by 9 makes the calculation easier

OpenStudy (anonymous):

can you help me on my other posts please